what is the differential equation having a solution y=cx²+cx?


1
Expert's answer
2021-09-12T23:52:33-0400

Consider the solution "y=cx+cx^2"

Require to find the differential equation from the given solution.

To find the differential equation, let us eliminate the parameter "c" by differentiating the given equation with respect to "x"

Differentiating the given equation "y=cx+cx^2" with respect to x , we get

"y'=c(1)+c(2x)"

Implies, we get "y'=c+2cx"

Implies, we get "y'=c(1+2x)"

"\Rightarrow c=\frac{y'}{1+2x}"

Substituting the value of "c" in the given equation "y=cx+cx^2" , we get

"\Rightarrow y=x[\frac{y'}{1+2x}]+x^2[\frac{y'}{1+2x}]"

"\Rightarrow y(1+2x)=xy'+x^2y'"

"\Rightarrow y(1+2x)=y'(x+x^2)"

Therefore, the required differential equation is

"y'=\frac{y(1+2x)}{(x+x^2)}"